A child is running on a moving sidewalk in an airport. When he runs against the sidewalk's motion, he travels
91ft in 13 seconds. When he runs with the sidewalk's motion, he travels
195ft in the same amount of time. What is the rate of the child on a still sidewalk and what is the rate of the moving sidewalk?

Answer by fractalier(6550)

Call the child's normal rate, r.
Call the rate of the walkway, w.
(r + w)(13) = 195
(r - w)(13) = 91
and
13r + 13w = 195
13r - 13w = 91
Now add these and get
26r = 286
r = 11 ft/s
By plugging in we find
w = 4 ft/s